OpenStudy (anonymous):

a. estimate the solution of this system of equations by graphing b. solve the system in part (a) using substation y=x-4 y= 1/3x-9

OpenStudy (anonymous):

What exactly are you having trouble with

OpenStudy (anonymous):

graphing and substituting

OpenStudy (zpupster):

what is your estimate??

OpenStudy (anonymous):

-11,-8

OpenStudy (zpupster):

lets use substitution and solve since y = x-4 x-4=1/3x-9 adding 4 to both sides x=1/3x-9+4 no we subtract 1/3x from both 3/3x-1/3x = -5 2/3x=-5 multiply by 3 and divide by 2 x=-15/2 or -7.5 back x-4=y -15/2-8/2=y -23/2=y or -11.5